### v1

@@ 4 Addressing and Mitigating Unfairness in AI
As AI technologies increasingly influence healthcare delivery, it becomes essential to scrutinize and refine these systems to prevent disparities in care and outcomes. This section explores various bias detecting and mitigating strategies in AI systems in healthcare.
@@ 4.1 Bias Detection Methods
In healthcare, detecting bias in AI systems is crucial, as it can result in disparities in healthcare delivery and outcomes. This section provides an overview of bias detection strategies in healthcare AI system, including: statistical analysis, auditing tools, and feedback from end-users.
4.1.1 Statistical Analysis:
In the realm of AI healthcare, detecting bias through statistical analysis is a critical step towards ensuring equitable healthcare outcomes across diverse populations. Regular statistical reviews serve as a foundational approach to identifying and addressing disparities in AI outputs, thereby improving the fairness and effectiveness of AI-driven healthcare solutions 123. The systematic review on bias detection and mitigation strategies in electronic health record-based models shows how important it is to find different types of bias that can affect AI models created using EHR data. This comprehensive analysis highlights the necessity of employing fairness metrics such as statistical parity 143, equal opportunity 144, and predictive equity 145 to detect implicit and algorithmic biases effectively. Using chi-squared statistical analysis, differences in preconception health indicators have been found based on age, race or ethnicity, education level, living in a city, and income. This shows how important it is to do regular statistical reviews to find healthcare inequalities 146. Similarly, a study of the clinical predictive values for radiographic pneumonia in children showed the sensitivity and specificity of commonly assessed signs and symptoms. This gave us information about how well AI models can be used for diagnosis in pediatric healthcare 147, 148. The epidemiological study of autism spectrum disorders in Greece used statistical analysis to find differences in the rates of ASD across different regions. This shows how useful statistical methods are for finding biases in healthcare datasets 149. Additionally, a prospective diagnostic evaluation comparing self-taken and healthcare worker-taken swabs for rapid COVID-19 testing highlighted the accuracy of AI-driven diagnostic tools, further illustrating the importance of statistical analysis in evaluating AI model performance 150. Furthermore, the exploration of AI in enhancing diagnosis, treatment, and healthcare systems in India emphasizes the potential of AI to revolutionize healthcare while also acknowledging the challenges posed by biases and ethical considerations 151. This underscores the need for rigorous statistical analysis to ensure that AI-driven healthcare solutions are both effective and equitable.
4.1.2 Auditing Tools
The Disparities Impact Statement (DIS): An innovative tool, the DIS 152, evaluates healthcare policies and practices for potential biases. It assesses how different population groups might be differently affected by a healthcare policy or practice, aiming to preemptively identify disparitiesṪhe tool involves qualitative and quantitative analyses to ensure a comprehensive review of the policy’s implications across various demographics 153. By implementing DIS, healthcare institutions can address disparities at the policy formulation stage, promoting equity in healthcare delivery 154.
PROGRESS Framework: The PROGRESS framework 155, which stands for Place of Residence, Race, Occupation, Gender, Religion, Education, Socioeconomic Status, and Social Capital, is a tool used to identify and evaluate the factors contributing to health inequalities. This framework encourages researchers and practitioners to systematically consider these dimensions when developing and implementing studies or interventions in healthcare settings 156. It is particularly useful in designing public health interventions and assessing the socio-economic and demographic factors influencing health outcomes 157.
Implicit Association Test (IAT): The IAT 158 is employed in healthcare to detect unconscious biases among healthcare providers that may affect patient care and treatment outcomes. Originally developed in psychological research to explore unconscious biases, IAT has been adapted for healthcare to assess biases in race, age, and obesity, among other aspects 159. Studies utilizing IAT have demonstrated that even well-intentioned staff can harbor biases that influence their decisions and interactions with patients 160.
Healthcare Disparities Dashboard: The Healthcare Disparities Dashboard 161 is a data-driven tool that aggregates patient data to highlight discrepancies in healthcare outcomes, utilization, and satisfaction across different demographic groups. It allows healthcare organizations to visualize and monitor disparities, thus aiding in targeted interventions 162. By presenting data in an accessible format, the Dashboard supports ongoing monitoring and the proactive management of healthcare equity 163.
Equity Quality Improvement (EQI) Tool: Healthcare systems design the EQI tool 164 to integrate equity into their quality improvement frameworks. This tool helps identify areas where healthcare disparities exist and suggests practical interventions to mitigate these disparities. The EQI tool facilitates the inclusion of equity as a central quality dimension, ensuring that improvements in healthcare delivery uniformly benefit all patient groups 165.
Cultural Competence Assessment (CCA): The CCA tool 166 helps healthcare institutions assess and enhance their effectiveness in providing culturally sensitive care. This tool evaluates healthcare providers’ awareness of and sensitivity to the cultural needs of diverse patient populations. The CCA can lead to improved patient satisfaction by fostering better communication and understanding between healthcare providers and patients from diverse backgrounds 167.
4.1.3 Feedback from End-Users
Collecting and analyzing feedback from healthcare providers who use AI systems is essential for identifying and mitigating biases that may affect clinical decisions and patient outcomes. Healthcare professionals are at the forefront of witnessing the operational aspects of AI, and their firsthand insights can reveal how these systems function across diverse clinical scenarios 5. Studies show that user feedback is invaluable for improving AI models, especially to ensure that these tools are equally effective and free from bias across different populations 168. By engaging healthcare providers in the evaluation process, developers can gather critical data on AI’s practical implications and areas where it may unintentionally perpetuate disparities 135. Moreover, incorporating such feedback facilitates the development of more robust, equitable, and contextually appropriate AI solutions, ultimately leading to better patient care and outcomes 4. As a result, systematic feedback mechanisms are not only beneficial but also necessary for the iterative refinement and ethical deployment of healthcare AI systems 137.
@@ 4.2.1 Diverse and Representative Data
Mitigating biases in AI systems within healthcare is crucial for ensuring equitable treatment across all patient demographics. A fundamental strategy involves using diverse and representative datasets in the training phase of these systems. Research indicates that AI models can only be as unbiased as the data they are trained on 169. Thus, inclusion of comprehensive data from a wide array of patient groups, particularly those historically underrepresented in medical research, is essential 13.
The significance of representative data extends beyond just including diverse groups; it also involves the depth and quality of the data collected from these groups. Data must capture a breadth of variables that influence health outcomes, such as socioeconomic factors, environmental conditions, and genetic differences 4. This approach helps in creating models that are better tuned to the nuances of various patient needs and conditions 14.
Additionally, involving stakeholders from diverse backgrounds in the data collection and AI development process can enhance the relevancy and sensitivity of the datasets 170. This inclusion helps in identifying and addressing potential blind spots in AI training datasets, which, if overlooked, could perpetuate biases and inequalities in healthcare delivery 171. Employing these strategies not only improves the fairness and effectiveness of AI tools but also builds trust in these technologies among all user groups, thereby fostering a more inclusive healthcare ecosystem 137.
@@ 4.2.2 Fairness-aware approaches
In healthcare, pre-processing involves techniques to adjust the input data for AI models to ensure it accurately represents diverse patient demographics before model training begins. This process aims to eliminate any inherent biases that may skew AI predictions and outcomes, thus fostering equity in healthcare treatments and diagnostics 172. By refining the dataset upfront, pre-processing helps in building AI systems that perform fairly across all patient groups 13.
  • Re-Sampling: Re-sampling is a common pre-processing method that manipulates data samples to create a more balanced dataset. This can be done through either over-sampling minority classes or under-sampling majority classes 173. An example of this in healthcare is the use of the Synthetic Minority Oversampling Technique (SMOTE) 174 in medical datasets where certain conditions or outcomes are rare. Chawla applied SMOTE to a dataset of imbalanced classes to improve the prediction of minority-class outcomes in medical diagnostics, resulting in more reliable predictive models for underrepresented conditions. This technique helps to ensure that the AI does not become biased towards the more frequent classes.
  • Re-Weighting: To reduce bias, instances in the training data are reweighted. We assign a weight to each instance in a dataset based on its representation, thereby increasing the influence of underrepresented groups in the model training process. This technique was demonstrated by Calmon in 175, who developed a data pre-processing method that modifies features and outcomes to improve fairness. In healthcare, this could mean adjusting weights so that data from minority ethnic groups has greater influence on the model, helping to prevent the AI from developing biases that might affect diagnosis or treatment recommendations 172.
@@
In-processing methods involve integrating fairness directly into the learning algorithm itself. This technique adjusts the algorithm during the training phase to minimize bias. Zemel in his study introduced a method where they developed a representation for data that is invariant to protected attributes like race or gender while maintaining the informative characteristics necessary for prediction 176. In healthcare, this can be crucial for ensuring that diagnostic or treatment recommendations are not adversely skewed by underlying biases in the training data 177.
  • Adversarial Debiasing: Adversarial debiasing involves the simultaneous training of a predictor and an adversary. The predictor learns to make accurate predictions, and the adversary learns to determine whether the predictions are biased toward certain groups. Deldjoo et al. 177 explored this in the context of healthcare, aiming to create clinical decision support systems that are unbiased towards patients’ demographic attributes. This technique helps ensure that the AI’s treatment recommendations do not reflect discriminatory patterns that might exist in the training data 178.
  • Constraint-Based Optimization: Constraint-based optimization involves modifying the learning algorithm to satisfy fairness constraints while minimizing prediction error. You can implement this method by adding fairness constraints to the objective function the algorithm is optimizing. Elzayn et al. 179 applied this method to healthcare datasets, ensuring that treatment recommendations do not disproportionately favor one group over another by enforcing demographic parity or equality of opportunity as constraints during model training 180.
@@
After training a model, post-processing techniques are applied to modify the outputs of AI systems, ensuring fair treatment across different demographic groups. These adjustments address any residual biases to promote equitable outcomes. For example, Hardt et al. 144 applied different decision thresholds for various groups to equalize treatment effects. This method effectively adjusted risk scores in a healthcare dataset, ensuring that treatment recommendations were equitable 181.
@@
  • Threshold Adjustment: Threshold adjustment in AI healthcare involves setting an optimal cutoff point for predictive models to classify outcomes effectively, such as predicting disease presence or patient risk levels. A significant example of this is in cardiovascular disease prediction. In a study by Siontis 182, the researchers examined the application of different thresholds for predicting cardiovascular events using a logistic regression model. They aimed to optimize the sensitivity and specificity of predictions by adjusting the threshold to better serve clinical decision-making. This adjustment proved crucial in identifying higher-risk patients who might benefit from preventive treatments. Another notable case is in breast cancer screening, where Fenton 183 applied threshold adjustment to mammography interpretation models. By altering the decision threshold, they were able to reduce false positives without substantially missing cases of cancer, thereby balancing the need for early detection with the risk of overdiagnosis and unnecessary anxiety for patients.
@@
  • Output Recalibration: Output recalibration in AI healthcare is critical for adapting predictive models to local contexts and ensuring that the predicted probabilities match actual clinical outcomes in a new patient population. A notable case involves the recalibration of a sepsis prediction model at Johns Hopkins Hospital, as detailed by Henry et al. 184. The model was originally developed with a dataset from one patient demographic, but recalibration using local demographic data was required to maintain accuracy across diverse populations within the hospital. This recalibration helped align the model’s predictions with the actual rates of sepsis observed, improving clinical decision-making and patient care. Another example is from the University of California, San Francisco (UCSF), where Rajkomar 4 recalibrated Google’s deep learning model used for predicting medical outcomes. The model, initially trained on a national dataset, was recalibrated with UCSF-specific patient data to ensure its predictions accurately reflected the local patient population’s risk factors and outcomes. This was particularly important for diseases with variable manifestations across different demographics.
5 Ethical and legal considerations
This section discusses the various ethical and legal considerations that are necessary to reduce bias in AI systems in healthcare. It will cover ethical issues such as privacy and data protection, bias in AI algorithms, transparency and explainability of AI decisions, potential displacement of healthcare professionals, and accountability for AI-driven decisions. Additionally, it will explore legal aspects including intellectual property rights, liability for AI-induced errors, compliance with healthcare regulations, and informed consent for AI involvement in patient care.
5.1 Ethical Considerations
The integration of AI in healthcare presents significant ethical considerations that necessitate careful scrutiny. One primary concern is privacy and data protection, as AI systems typically require large datasets, including sensitive patient information, to train and operate effectively. Ensuring the confidentiality and security of this data against breaches is paramount 185. Additionally, the issue of bias in AI algorithms poses a serious ethical challenge. AI systems can perpetuate and even exacerbate existing biases if they are trained on non-representative data, leading to unequal treatment outcomes among different demographic groups 4. There is also the ethical issue of transparency and explainability. Practitioners and patients may find it challenging to comprehend the reasoning behind AI decision-making processes due to their opaque nature. This lack of transparency can undermine trust in AI systems and complicate informed consent processes 186. Furthermore, the deployment of AI in healthcare must take into account the potential displacement of healthcare professionals, raising concerns about job security and the loss of valuable human expertise and empathy in patient care 187. Lastly, accountability for AI-driven decisions in healthcare is a critical ethical issue. It is critical to define who is responsible—the developers, the users, or the AI itself—when an AI system’s decision results in patient harm 188.
5.2 Legal Considerations
The deployment of AI in healthcare brings forth numerous legal considerations that are crucial for ensuring compliance and safeguarding patient rights. Intellectual property rights stand out prominently, as there is ongoing debate over whether AI-generated medical inventions can be patented and who owns the rights to AI-generated data and algorithms 138. Furthermore, liability issues arise when AI systems make erroneous decisions that could harm patients. Determining whether the healthcare provider, the software developer, or another party is liable requires intricate legal analysis and possibly new legal frameworks 189. Compliance with medical standards and regulatory requirements is another critical legal area. AI applications in healthcare must adhere to existing healthcare regulations, such as the Health Insurance Portability and Accountability Act (HIPAA) in the U.S., which governs the privacy and security of patient data, and the General Data Protection Regulation (GDPR) in the EU, which sets stringent guidelines for data protection and privacy 190. As AI technologies evolve, there may also be a need for specific regulations that address AI’s unique aspects in healthcare to ensure that these innovations are safely and effectively integrated into medical practice 191. Lastly, issues of informed consent are magnified with AI, as patients must understand the role of AI in their care, including how their data will be used and the implications of AI-driven decisions 188.

### v2

@@ 3 Addressing and Mitigating Unfairness in AI
As AI technologies increasingly influence healthcare delivery, it becomes essential to scrutinize and refine these systems to prevent disparities in care and outcomes. This section explores various bias-detecting and mitigating strategies in AI systems in healthcare.
@@ 3.1 Bias Detection Methods
Bias detection is a crucial step in identifying and addressing inequities in AI systems, especially in healthcare given its high-stakes nature. This section discusses various tools and techniques for detecting bias, categorized by their roles in the pre-processing, in-processing, and post-processing stages of the machine learning (ML) pipeline. The aim is to provide a clear and actionable understanding of bias detection.
3.1.1 Pre-Processing Stage
The pre-processing stage focuses on detecting biases in datasets before model training begins. Identifying these biases at this stage is essential to prevent their propagation and amplification through the ML pipeline if left unchecked. Specifically, bias can be detected using Fairness Metrics Techniques, such as Equalized Odds and Equal Opportunity Difference, which measure disparities in clinical outcomes or feature distributions across demographic groups. These metrics help identify biases in healthcare datasets, such as selection bias and measurement bias by revealing whether certain patient populaitons are disadvantaged or overrepresented. For instance, if one group consistently experiences worse outcomes in the data, these tools can pinpoint the issue, enabling targeted corrections [88].
Another detection method is Dataset Visualization Techniques, such as t-SNE (t-Distributed Stochastic Neighbor Embedding) and PCA (Principal Component Analysis), which allow researchers to examine bias in high-dimensional healthcare data in more intuitive ways. These tools uncover patterns, such as clusters or representation gaps, that may signal clinical representation bias. For example, a t-SNE plot could reveal that certain skin tones are underrepresented in a dermatological dataset, prompting actions to balance the dataset [94]. Another approach is the use of Causal Graphs, which focus on identifying and understanding the relationships between variables in healthcare datasets. Causal Graphs are graphical representations of causal relationships among variables, allowing researchers to distinguish true causation from correlations and uncover confounding variables that may introduce bias. For instance, [95] proposed a framework leveraging Causal Variational Autoencoders (CVAEs) to indirectly reconstruct sensitive information, even when such attributes are unavailable due to privacy constraints. By identifying clinical biases at the data level, this method proves particularly valuable in addressing challenges within complex, real-world medical datasets.
3.1.2 In-Processing Stage
The in-processing stage focuses on detecting biases introduced or amplified during model training. This stage prioritizes fairness by incorporating bias identification strategies directly into the training process. One effective method is Adversarial Learning, which employs adversarial techniques to identify bias in models during training. A primary model predicts clinical outcomes (e.g., disease risk ), while an adversary attempts to infer sensitive attributes such as patient demographics (e.g., race, gender) based on the model’s predictions. The adversary’s success in predicting these attributes reveals the extent of algorithmic bias embedded in the model. By analyzing this feedback, healthcare developers can identify where and how protected attributes influence clinical predictions, enabling precise bias detection and promoting equitable care [96].
Another approach involves Explanation Methods, such as LIME (Local Interpretable Model-Agnostic Explanations), which provide insights into how clinical features influence predictions in healthcare models. These methods can uncover both explicit and implicit biases in the model, such as disproportionately relying on sensitive attributes like race, gender, or socioeconomic status or unintentionally reinforcing biased patterns that disadvantage certain patient groups. By iteratively applying these explanations during training, healthcare practitioners can identify and correct biases, improving fairness and transparency in patient care [97]. A further technique for bias detection is Causal Modeling, which identifies algorithmic bias by analyzing how sensitive attributes, such as gender, ethnicity, or socioeconomic status, influence predictive outcomes in healthcare applications. Causal Modeling involves constructing mathematical or computational models that map out causal relationships between variables, enabling researchers to trace pathways where biases emerge during the training of models used in medical decision-making. For instance, researchers utilized causal modeling with the Adult dataset from the UC Irvine Machine Learning Repository to detect gender bias in binary classification outcomes [98]. By treating the prediction model as a black box and constructing a corresponding causal model, they identified how gender impacted the model’s decisions. This process uncovers both explicit and implicit biases in the model’s decision-making framework, providing actionable insights for detecting inequities. By identifying biases at the model level, causal modeling promotes the development of fair and robust AI systems in health care.
3.1.3 Post-Processing Stage
The post-processing stage evaluates and mitigates any biases in the model’s outputs after training is complete. This stage ensures equitable treatment in the model’s final decisions. One effective approach for bias detection is Counterfactual Analysis, which assesses whether a model’s decisions remain consistent even if sensitive attributes, such as race or gender, are changed. By using causal inference, this method determines whether changing a sensitive attribute would alter the model’s outcome. If the outcome remains the same, the decision is considered fair, making this method effective for identifying and correcting both explicit and implicit biases in the model’s outputs [99] For instance, in a healthcare model that predicts eligibility for advanced treatments, if changing a sensitive attribute like race results in different treatment outcomes, it highlights bias in the model’s decision-making process.
Another tool is Model Cards, which are standardized documents that summarize a model’s performance across various clinical conditions and demographic groups. They include details about the model’s intended use in healthcare settings, limitations, and evaluation results, including the detection and analysis of evaluation bias. By promoting transparency, Model Cards help healthcare stakeholders make informed decisions about deployment and encourage responsible use [100]. Another detection approach, Reject Option involves identifying cases near the decision boundary where classification decisions are uncertain and potentially influenced by bias. This approach helps detect patterns where sensitive demographic groups might be disproportionately affected by misclassifications. In healthcare settings, where errors in diagnosis or treatment can have significant consequences, the reject option highlights instances that require further inspection to ensure equitable outcomes. By focusing on borderline cases, this method provides valuable insights into potential biases in model predictions, aiding in the identification of inequities within healthcare systems [101].
This framework provides a clear and actionable approach to tackling biases in AI systems. Detecting and identifying biases at each stage of the ML pipeline is essential for creating fair, reliable, and equitable AI-driven healthcare solutions.
@@ 3.2.1 Diverse and Representative Data
Mitigating biases in AI systems within healthcare is crucial for ensuring equitable treatment across all patient demographics. A fundamental strategy involves using diverse and representative datasets in the training phase of these systems. Research indicates that AI models can only be as unbiased as the data they are trained on [102]. Thus, the inclusion of comprehensive data from a wide array of patient groups, particularly those historically underrepresented in medical research, is essential [13].
The significance of representative data extends beyond just including diverse groups; it also involves the depth and quality of the data collected from these groups. Data must capture a breadth of variables that influence health outcomes, such as socioeconomic factors, environmental conditions, and genetic differences [4]. This approach helps in creating models that are better tuned to the nuances of various patient needs and conditions [14].
Additionally, involving stakeholders from diverse backgrounds in the data collection and AI development process can enhance the relevancy and sensitivity of the datasets [70]. This inclusion helps in identifying and addressing potential blind spots in AI training datasets, which, if overlooked, could perpetuate biases and inequalities in healthcare delivery [103]. Employing these strategies not only improves the fairness and effectiveness of AI tools but also builds trust in these technologies among all user groups, thereby fostering a more inclusive healthcare ecosystem [81].
@@ 3.2.2 Fairness-aware approaches
In healthcare, pre-processing involves techniques to adjust the input data for AI models to ensure it accurately represents diverse patient demographics before model training begins. This process aims to eliminate any inherent biases that may skew AI predictions and outcomes, thus fostering equity in healthcare treatments and diagnostics [104]. By refining the dataset upfront, pre-processing helps in building AI systems that perform fairly across all patient groups [13].
  • Re-Sampling: Re-sampling is a common pre-processing method that manipulates data samples to create a more balanced dataset. This can be done through either over-sampling minority classes or under-sampling majority classes [105]. An example of this in healthcare is the use of the Synthetic Minority Oversampling Technique (SMOTE) [106] in medical datasets where certain conditions or outcomes are rare. Chawla applied SMOTE to a dataset of imbalanced classes to improve the prediction of minority-class outcomes in medical diagnostics, resulting in more reliable predictive models for underrepresented conditions. This technique helps to ensure that the AI does not become biased towards the more frequent classes.
  • Re-Weighting: To reduce bias, instances in the training data are reweighted. We assign a weight to each instance in a dataset based on its representation, thereby increasing the influence of underrepresented groups in the model training process. This technique was demonstrated by Calmon in [107], who developed a data pre-processing method that modifies features and outcomes to improve fairness. In healthcare, this could mean adjusting weights so that data from minority ethnic groups has a greater influence on the model, helping to prevent the AI from developing biases that might affect diagnosis or treatment recommendations [104].
@@
In-processing methods involve integrating fairness directly into the learning algorithm itself. This technique adjusts the algorithm during the training phase to minimize bias. Zemel in his study introduced a method where they developed a representation for data that is invariant to protected attributes like race or gender while maintaining the informative characteristics necessary for prediction [108]. In healthcare, this can be crucial for ensuring that diagnostic or treatment recommendations are not adversely skewed by underlying biases in the training data [109].
  • Adversarial Debiasing: Adversarial debiasing involves the simultaneous training of a predictor and an adversary. The predictor learns to make accurate predictions, and the adversary learns to determine whether the predictions are biased toward certain groups. Deldjoo et al. [109] explored this in the context of healthcare, aiming to create clinical decision support systems that are unbiased towards patients’ demographic attributes. This technique helps ensure that the AI’s treatment recommendations do not reflect discriminatory patterns that might exist in the training data [110].
  • Constraint-Based Optimization: Constraint-based optimization involves modifying the learning algorithm to satisfy fairness constraints while minimizing prediction error. You can implement this method by adding fairness constraints to the objective function that the algorithm is optimizing. Elzayn et al. [111] applied this method to healthcare datasets, ensuring that treatment recommendations do not disproportionately favor one group over another by enforcing demographic parity or equality of opportunity as constraints during model training [112].
@@
After training a model, post-processing techniques are applied to modify the outputs of AI systems, ensuring fair treatment across different demographic groups. These adjustments address any residual biases to promote equitable outcomes. For example, Hardt et al. [88] applied different decision thresholds for various groups to equalize treatment effects. This method effectively adjusted risk scores in a healthcare dataset, ensuring that treatment recommendations were equitable [113].
@@
  • Threshold Adjustment: Threshold adjustment in AI healthcare involves setting an optimal cutoff point for predictive models to classify outcomes effectively, such as predicting disease presence or patient risk levels. A significant example of this is in cardiovascular disease prediction. In a study by Siontis [114], the researchers examined the application of different thresholds for predicting cardiovascular events using a logistic regression model. They aimed to optimize the sensitivity and specificity of predictions by adjusting the threshold to better serve clinical decision-making. This adjustment proved crucial in identifying higher-risk patients who might benefit from preventive treatments. Another notable case is in breast cancer screening, where Fenton [115] applied threshold adjustment to mammography interpretation models. By altering the decision threshold, they were able to reduce false positives without substantially missing cases of cancer, thereby balancing the need for early detection with the risk of overdiagnosis and unnecessary anxiety for patients.
@@
  • Output Recalibration: Output recalibration in AI healthcare is critical for adapting predictive models to local contexts and ensuring that the predicted probabilities match actual clinical outcomes in a new patient population. A notable case involves the recalibration of a sepsis prediction model at Johns Hopkins Hospital, as detailed by Henry et al. [116]. The model was originally developed with a dataset from one patient demographic, but recalibration using local demographic data was required to maintain accuracy across diverse populations within the hospital. This recalibration helped align the model’s predictions with the actual rates of sepsis observed, improving clinical decision-making and patient care. Another example is from the University of California, San Francisco (UCSF), where Rajkomar [4] recalibrated Google’s deep learning model used for predicting medical outcomes. The model, initially trained on a national dataset, was recalibrated with UCSF-specific patient data to ensure its predictions accurately reflected the local patient population’s risk factors and outcomes. This was particularly important for diseases with variable manifestations across different demographics.